# Blueberry galaxies up to 200 Mpc and their optical/infrared properties

### v1

@@ 3 Results
MPA refers to results provided by MPA-JHU, HEC to HECATE, S to spectral fitting, and SED to SED fitting.
Distances are as reported in HECATE except for objects with spectroscopic observations where distances are based on their redshift.
Y=Yes, and N=No if the BB is compact based on visual inspection.
@@ 3.2 Sizes and morphology
The size distribution of the selected BBs is shown in Figure 3. The sizes were estimated based on the major axes reported in HECATE, and distances of the galaxies as reported in Table 3 The mode of their size distribution is kpc. Visual inspection of this work’s BBs (Figure 2) shows that the majority of BBs appear to be compact and isolated sources with a spherical morphology. However, some sources (BBs 1, 2, 3, 5, 6, 13, 15, 16, 20, 24, 31, 33, 34, 37, 39, 40, 42, 43, 45, 46, 48) show extended morphology. Some of these BBs may be blue star-forming regions that are part of a slightly larger galaxy where a part of it has less SFR in comparison (e.g., BBs 3, 5, 6, 33, 34, 42). For instance, this is indeed the case for BB 5 which has spectroscopic observations for the bluest ( Å; north-east), and the less-intensive ( Å; south-west) regions of the source, both of which share the same redshift. The optical and IR photometries correspond to the full extent of the galaxies, thus, these sources fulfill the selection criteria even when accounting for their less intensive regions. In some cases, nearby angular-projection sources may be background, more distant, and unrelated galaxies (e.g., possibly BB 46). Spectroscopic observations at different locations on the surface of these extended features would be required to identify whether these regions belong to these galaxies.
The mode and C.Is. of the distribution of distances for the compact BBs is (Mpc), and for the ones with extended features is (Mpc). Therefore, we do not find a strong dependence on whether the BBs appear compact due to the combination of large distances and their small overall size. However, the two nearest sources show extended morphology (BBs 3, and 42), Moreover, some recent studies have shown that some spatially resolved GP analogs show irregular morphology (e.g., Mrk 71 at 3.4 Mpc; Micheva et al. 2017) while deep HST observations also revealed extended structures in GPs (e.g., Jaskot et al. 2019).
The three largest reported BBs (6, 34, and 24) show extended features. However, among the largest are BBs 22 and 41, both having a size of 5.4 kpc, which appear as compact sources. However, these are exceptional sources as the mode of the size distribution of compact BBs is kpc. A unique source is BB 43 which appears to be a binary BB galaxy comprised of two compact sources that otherwise seem to be isolated from other nearby galaxies. Unfortunately, there are no spectra available for BB 43 while its photometry accounts for both sources.
@@ 3.3 Spectral energy distribution fitting
To homogeneously estimate stellar masses and SFRs for all the reported BBs, we performed spectral energy distribution (SED) fitting taking advantage of the multiwavelength photometric coverage of the sources. In addition to the visible SDSS, Pan-STARRS, and IR WISE the sources were cross-matched with the GALEX extended source catalog (Martin et al. 2005) in order to include UV photometry. The Blueberry galaxies BB 1, 3, 38, 42, 45, and 46 were not detected by GALEX in any of the AIS, GII, or MIS surveys.
@@
The SED fitting was performed with the updated version (Yang et al. 2022) of the widely used code CIGALE (Burgarella et al. 2005; Noll et al. 2009; Boquien et al. 2019). Particularly, we used the sfhdelayed, bc03, nebular, dustatt_modified_starburst, redshifting, and dale2014 modules covering a wide range of the parameter space. We focused the parameter grid on low-metallicity, and high sSFR galaxies, which is the case of BBs. For instance, the sfhdelayed module f_burst parameter which refers to the relative strength of the recent star-formation burst, was allowed to be even up to 0.99 and the age_burst down to 1 Myr. We also allowed for a 10% systematic uncertainty in the observed fluxes. All sources were fitted with components only related to star formation. Overall, the applied grid resulted in 5,225,472 models for each BB galaxy, and the best fit was chosen based on the minimum reduced- (). All sources had good fits with average and highest for BB 20. However, our SFR and estimations were not based on the best-fit model but on their -weighted probability density functions (PDFs). Figure 12 shows the SED plots of BB 20 and BB 30 as indicative of the SED fitting process.
@@
Appendix B provides the -weighted PDFs of the SFR and for each galaxy separately. These PDFs cover all the models created by CIGALE in the fitting process, and thus, this approach allows us to evaluate how much the SFR and estimations vary for the different models and obtain a more robust estimation of the parameter quantities and uncertainties. The quantity/error estimations provided by Yang et al. (2022) correspond to the average/standard deviation of these PDFs. However, because in our case most of the PDFs do not have a Gaussian form, we adopt and calculate the estimations based on the mode and 68% C.Is.
@@
Since MPA-JHU also provides stellar masses and SFRs, we compare these estimations with our SED fitting for the 21 BBs of this work analyzed by MPA-JHU. The logarithm of the ratio of the SED fitting and MPA-JHU SFRs is . Similarly, compared to the HECATE SFRs . Thus, the SFR estimations of the SED fitting tend to give larger SFRs. The logarithmic ratio of the SED fitting and MPA-JHU stellar masses is , and . Thus, the stellar masses derived by the SED fitting tend to agree more with MPA-JHU and less with HECATE, however, there is considerable scatter.
@@
The different methods presented here for the SFR and estimations may suffer from the same or other biases. Galaxies like the BBs may not be very well represented by the CIGALE assumption of a standard H II cloud with electron density and electron temperature K. For instance, the nearest GP analog Mrk 71 shows and K (Micheva et al. 2017). On the other hand, MPA-JHU SFRs were estimated through a combination of the emission line measurements from the fiber spectra, with aperture corrections that were applied with the method of Gallazzi et al. (2005), and the photometric measurements outside the fiber using the method of Salim et al. (2007). Only integrated photometry was used for sources with no adequate SFR estimation from the emission lines. The SFRs and stellar masses, in HECATEv2.0 are based on IR WISE photometry and combinations with optical photometry or estimations of extinction depending on availability (Kouroumpatzakis et al. 2023). It is possible that in some cases MPA-JHU may have underestimated the SFR of the BBs since they did not include UV or mid-IR emission and may not have fully accounted for extinction. Both are important given that BBs show strong UV emission. The IR-based estimations of the HECATE SFRs may underestimate the SFRs of high-sSFR and low-extinction galaxies like the BBs for the same reason (e.g., Kouroumpatzakis et al. 2023).
@@ 3.4 Blueberry galaxies with respect to the main sequence of star-forming galaxies
To place the proposed BBs in the context of their current star-forming conditions we plot their SFR as a function of , compared to the main sequence of SFGs. Figure 4 shows the proposed BBs, and the samples of Cardamone et al. (2009), Izotov et al. (2011), Yang et al. (2017), and all spectroscopically classified SFGs (class_sp = 0) of the HECATE catalog. The SFRs and stellar masses plotted here refer to the results of the SED fitting for the BBs of this work (Section 3.3). For the Yang et al. (2017) BBs, SFRs, and stellar masses are as reported in their work. For the HECATE catalog, and the samples of Cardamone et al. (2009), and Izotov et al. (2011), the SFRs and stellar masses are as reported in HECATEv2.0, and the MPA-JHU catalogs respectively.
All aforementioned samples of dwarf galaxies show very high values of sSFR () and are placed above the main sequence compared to HECATE galaxies which represent here the average population of SFGs in the near Universe. Moreover, BBs tend to be even more extreme compared to the Cardamone et al. (2009) GPs and Izotov et al. (2011) BDGs, with the majority being less massive and having .
@@ 3.5 Visible colors of Blueberry galaxies
Figure 5 shows the rest-frame optical color-color versus diagram for this work’s BBs, the Yang et al. (2017) BBs, the Izotov et al. (2011) BDGs, and the Cardamone et al. (2009) GPs. The HECATE galaxies classified as star-forming are also shown, representing the general population of SFGs in the nearby Universe. The colors plotted for both BB samples and the HECATE SFGs are based on SDSS cModelMag photometry. The colors of the BDGs and GPs are based on ModelMag photometry from SDSS DR12 (Alam et al. 2015). The photometry of all sources was -corrected using the conversions of Chilingarian et al. (2010) and Chilingarian & Zolotukhin (2012). Figure 5 shows that some BBs are redder than , which is the selection threshold adopted by this work for SDSS photometry. However, these sources were selected through Pan-STARRS photometry which, as discussed in Section 2, can be different compared to the cModelMag plotted here (see also Table 2). Colors reported in Table 2 refer to the observed, not the -corrected colors.
@@
The majority of this work’s BBs tend to be slightly redder in the color compared to Yang et al. (2017) BBs but extend similarly in the color. However, there are a few sources that are located at the blue-end part of the color extending to the Yang et al. (2017) BBs. Overall, this work’s BBs tend to be located in the vs diagram between the Yang et al. (2017) BBs, and the Izotov et al. (2011) BDGs and Cardamone et al. (2009) GPs. This comparison also shows that the GPs and BDGs are on average intrinsically redder compared to BBs, although a small fraction of the large sample of Izotov et al. (2011) BDGs share similar blue colors with the BBs. Thus, the distinctive visible colors of BBs and GPs are not only different due to redshifting but are intrinsically different.
3.6 Spectroscopic properties of Blueberry galaxies.
@@
To examine their spectroscopic properties, the BBs were cross-matched with the MPA-JHU catalog which provides robust results based on a homogeneous spectroscopic analysis. However, this analysis can only be performed for the sources covered by the MPA-JHU catalog, that is 21/48 galaxies. The EWs of the , H, , and H emission lines, and the metallicities for the spectroscopic subsample of the proposed BBs are reported in Table 3 based on the MPA-JHU catalog. The distributions of the sSFR, and the EWs of the H, , and H emission lines for the GP and BB samples are presented in Figure 6. The spectroscopically classified HECATE SFGs are also shown. Emission-line fluxes and EWs for the Cardamone et al. (2009), Izotov et al. (2011), HECATE, and BBs of this work are as provided by the MPA-JHU analysis. Spectroscopic information of the Yang et al. (2017) BBs is as reported in their work but because many of these sources were observed with MMT, with poor flat-fielding of the red part of the spectrum, their were not properly estimated and are not shown in Figure 6. Table 4 presents the modes and 68% C.Is. of these distributions summarizing the results shown in Figure 6.
@@
The Blueberry galaxies of this work are in agreement and show similar sSFR distribution with the BBs of Yang et al. (2017). Both samples are on the right end of the sSFR distributions showing the highest star-forming activity among the presented samples. While the Cardamone et al. (2009) GPs and Izotov et al. (2011) BDGs also show very high sSFRs () compared to the main population of SFGs, BBs show even higher star-formation intensity. The comparison of the sample’s EW distributions shows a similar behavior to the one of the sSFRs. This work’s BBs show a similar EW distribution to the Yang et al. (2017) BBs. Both samples of BBs have the highest EWs in any of the compared emission lines.
@@
The sSFR of this work’s BBs mentioned in this table refers to the MPA-JHU estimations regarding the 21 spectroscopic sources. The modes and C.Is. of the sSFR of all 48 BBs based on the SED fitting is .
@@
The distributions of log for the different samples are shown in Figure 6. The ratio of the different oxygen ions is a good tracer of the interstellar medium’s (ISM) ionization degree. The distributions of GPs and BDGs overlap to some extent with the BBs, but the modes of their distributions have lower ratios, thus, BBs show the most ionized ISM among the compared samples.
@@
The mass–metallicity relation of BBs, GPs, and BDGs is presented in Figure 7. Metallicities and stellar masses of all samples, except the Yang et al. (2017) BBs, are as reported by the MPA-JHU analysis, where the metallicity estimation was based on the Charlot & Longhetti (2001) models. Stellar masses and metallicities of the Yang et al. (2017) BBs are as reported in their work. Most BBs show extremely low metallicities with . The mode of the metallicity distribution of the Yang et al. (2017) BBs is , and of this work’s BBs . For reference, we also plot the mass–SFR–metallicity relation by Curti et al. (2020) provided in their Equation 2. Estimations below are an extrapolation since the Curti et al. (2020) relation was a product of an analysis with more massive galaxies compared to BBs.
@@
The highly star-forming and low-mass BBs tend to have metallicities close to what is expected from the relation accounting for their stellar mass and SFR. For instance, the mode of the SFR distribution of the plotted BBs of this work is while the minimum and maximum SFRs are , respectively. The upper line drawn from the Curti et al. (2020) relation refers to , thus the majority of BBs are below the lines that correspond to their actual SFRs. However, the comparison shows significant scatter. GPs and most BDGs are more massive galaxies and show on average higher metallicities compared to BBs.
@@
There are also two outliers showing higher metallicity compared to the rest of the BBs (BBs 2 and 33). Both of these sources show extended features, rather than the typical compact morphology of the majority of BBs, and their spectroscopic analysis refers to these regions. This suggests that these extended features are part of the same galaxy which probably hosts older SPs and has ISM with higher metallicity.
@@
The comparison between the metallicity and the extinction of the aforementioned samples is shown in Figure 8. The extinction here is based on the Balmer ratio and the emission-line fluxes measured by MPA-JHU for the spectroscopic subsample of the proposed BBs. The was calculated by adopting Case B recombination with (Osterbrock 1989), and the reddening law of Calzetti et al. (2000). The of sources that were found to have a Balmer ratio was set to zero. The respective extinction–metallicity relation by Kouroumpatzakis et al. (2021), which was based on a representative sample of all types of SFGs, is also drawn. Figure 8 does not show the Yang et al. (2017) sample because for some galaxies there were problems with the red part of their spectrum. While the majority of the Cardamone et al. (2009) GPs, and Izotov et al. (2011) BDGs tend to follow the relation, given the large scatter, this is not the case for the majority of BBs. Based on their metallicity, most of this work’s BBs are below the relation and show stronger extinction than expected.
@@
Table 4 also provides the modes and 68% C.Is. of the distributions of the quantities discussed above, separately for the morphologically compact BBs and those showing extended features. On average, the compact sources tend to form stars more actively showing slightly higher sSFR and EWs. On the other hand, compact BBs seem to have on average higher metallicities although their distributions overlap and the BB with the highest metallicity (BB 2; 12 + log(O/H) = 8.6) shows extended features. Out of the 21 spectroscopically observed BBs, only 6 sources are reported as extended and 15 as compact, thus, these results should be taken with caution due to the small and uneven size of these sub-samples.
@@
The high ionization potential of (54.42 eV) is considered in many cases an indication of AGN activity. However, when AGN is not responsible, such high ionization levels can only be achieved by the presence of very massive stars and recent bursts of star-formation that generate SPs with ages less than 5 Myr (e.g., Bergeron et al. 1997). The emission-line fluxes were measured by Shirazi & Brinchmann (2012) for the spectra of SDSS Data Release 7. In total, 8 BBs of this work were reported by Shirazi & Brinchmann (2012) as sources with detected emission with S/N.
@@
To homogeneously estimate the fluxes for all the BBs with spectroscopic observations we fit their SDSS spectra with the SP fitting code FADO (Gomes & Papaderos 2017). FADO fits were based on the Bruzual & Charlot (2003) simple SP template library with metallicities and . All fits used the spectral range Å while we adopted the recommended resolution for SDSS spectra of 2.3 Å and recessional velocity margins of 500 . The fluxes and EWs of all the emission lines estimated by FADO are provided in the online Table.
@@
A comparison between the flux estimations of our analysis and MPA-JHU shows excellent agreement with , , and . The comparison of the EW estimations is , , and . Thus FADO tends to yield slightly larger EWs on average, mainly for the Balmer lines. Out of the 21 sources with spectroscopic observations, 15 were found to have emission with showing that BBs are indeed dominated by extremely young SPs. These are the BBs 2, 7, 9, 10, 11, 12, 14, 18, 20, 22, 27, 30, 31, 32, 33, and 36.

### v3

@@ 3.1 Proposed sample of blueberry galaxies
”MPA” refers to results provided by MPA-JHU, ”HEC” to HECATE, ”S” to spectral fitting, and ”SED” to SED fitting.
Distances are as reported in HECATE except for objects with spectroscopic observations, where distances are based on their redshift.
Y=Yes and N=No if the BB is compact based on visual inspection.
@@ 3.2 Sizes and morphology
The size distribution of the selected BBs is shown in Figure 3. The sizes were estimated based on the major axes reported in HECATE and the distances of the galaxies as reported in Table 3. The mode of their size distribution is kpc. Visual inspection of this work’s BBs (Figure 2) shows that the majority of BBs appear to be compact and isolated sources with a spherical morphology. However, some sources (BBs 1, 2, 3, 5, 6, 13, 15, 16, 20, 24, 31, 33, 34, 37, 39, 40, 42, 43, 45, 46, 48) show an extended morphology. Some of these BBs may be blue star-forming regions that are part of a slightly larger galaxy, where a part of it has less SFR in comparison (e.g., BBs 3, 5, 6, 33, 34, 42). For instance, this is indeed the case for BB 5, which has spectroscopic observations for the bluest ( Å; northeast) and the less intensive ( Å; southwest) regions of the source, both of which share the same redshift. The optical and IR photometries correspond to the full extent of the galaxies. Thus, these sources fulfill the selection criteria even when accounting for their less intensive regions. In some cases, nearby angular-projection sources may be background and thus unrelated galaxies (e.g., possibly BB 46). Spectroscopic observations at different locations of the surface of these extended features would be required to identify whether these regions belong to these galaxies.
The mode and CIs of the distribution of distances for the compact BBs is (Mpc), and for the ones with extended features, it is (Mpc). Therefore, we did not find a strong dependence on whether the BBs appear compact due to the combination of large distances and their small overall size. However, the two nearest sources show an extended morphology (BBs 3 and 42), Moreover, some recent studies have shown that some spatially resolved GP analogs show an irregular morphology (e.g., Mrk 71 at 3.4 Mpc; Micheva et al. 2017), while deep HST observations have also revealed extended structures in GPs (e.g., Jaskot et al. 2019).
The three largest reported BBs (6, 34, and 24) show extended features. Among the largest are also BBs 22 and 41 (size = 5.4 kpc) which appear to be compact sources. However, these are exceptional sources, as the mode of the size distribution of compact BBs is kpc. A unique source is BB 43, which appears to be a binary BB galaxy comprised of two compact sources that otherwise seem to be isolated from other nearby galaxies. Unfortunately, there are no spectra available for BB 43, while its photometry accounts for both sources.
@@ 3.3 Spectral energy distribution fitting
To homogeneously estimate stellar masses and SFRs for all the reported BBs, we performed spectral energy distribution (SED) fitting, taking advantage of the multiwavelength photometric coverage of the sources. In addition to the visible SDSS, Pan-STARRS, and IR WISE, the sources were cross-matched with the GALEX extended source catalog (Martin et al. 2005) in order to include UV photometry. The galaxies BB 1, 3, 38, 42, 45, and 46 were not detected by GALEX in any of the AIS, GII, or MIS surveys.
@@
The SED fitting was performed with the updated version (Yang et al. 2022) of the widely used code CIGALE (Burgarella et al. 2005; Noll et al. 2009; Boquien et al. 2019). In particular, we used the sfhdelayed, bc03, nebular, dustatt_modified_starburst, redshifting, and dale2014 modules covering a wide range of the parameter space. We focused the parameter grid on low-metallicity and high sSFR galaxies, which is the case of BBs. For instance, the sfhdelayed module f_burst parameter, which refers to the relative strength of the recent star-formation burst, was allowed up to 0.99 and the age_burst was allowed down to 1 Myr. We also allowed for a 10% systematic uncertainty in the observed fluxes. All sources were fitted with components only related to star formation. Overall, the applied grid resulted in 5,225,472 models for each BB galaxy, and the best fit was chosen based on the minimum reduced (). All sources had good fits with the average and highest for BB 20. However, our SFR and estimations were not based on the best-fit model but on their -weighted probability density functions (PDFs). Figure 12 shows the SED plots of BB 20 and BB 30 as indicative of the SED fitting process.
@@
Appendix B provides the -weighted PDFs of the SFR and for each galaxy separately. These PDFs cover all the models created by CIGALE in the fitting process, and thus, this approach allowed us to evaluate how much the SFR and estimations vary for the different models and obtain a more robust estimation of the parameter quantities and uncertainties. The quantity and error estimations provided by Yang et al. (2022) correspond to the standard deviation of these PDFs. However, because in our case most of the PDFs do not have a Gaussian form, we adopted and calculated the estimations based on the mode and 68% CIs.
@@
Since MPA-JHU also provides stellar masses and SFRs, we compared these estimations with our SED fitting for the 21 BBs of this work analyzed by MPA-JHU. The logarithm of the ratio of the SED fitting and MPA-JHU SFRs is . Similarly, compared to the HECATE SFRs, . Thus, the SFR estimations of the SED fitting tend to give larger SFRs. The logarithmic ratio of the SED fitting and MPA-JHU stellar masses is and . Thus, the stellar masses derived by the SED fitting tend to agree more with MPA-JHU and less with HECATE; however, there is considerable scatter.
@@
The different methods presented here for the SFR and estimations may suffer from the same or other biases. Galaxies such as the BBs may not be very well represented by the CIGALE assumption of a standard H II cloud with electron density and electron temperature K. For instance, the nearest GP analog Mrk 71 shows and K (Micheva et al. 2017). On the other hand, MPA-JHU SFRs were estimated through a combination of the emission-line measurements from the fiber spectra, with aperture corrections that were applied with the method of Gallazzi et al. (2005), and the photometric measurements outside the fiber using the method of Salim et al. (2007). Only integrated photometry was used for sources with no adequate SFR estimation from the emission lines. The SFRs and stellar masses in HECATEv2.0 are based on IR WISE photometry and combinations with optical photometry or estimations of extinction, depending on availability (Kouroumpatzakis et al. 2023). It is possible that in some cases MPA-JHU may have underestimated the SFR of the BBs since they did not include UV or mid-IR emission and may not have fully accounted for extinction. Both are important given that BBs show strong UV emission. The IR-based estimations of the HECATE SFRs may underestimate the SFRs of high-sSFR and low-extinction galaxies such as the BBs for the same reason (e.g., Kouroumpatzakis et al. 2023).
@@ 3.4 Blueberry galaxies with respect to the main sequence of star-forming galaxies
To place the proposed BBs in the context of their current star-forming conditions, we plotted their SFR as a function of , compared to the main sequence of SFGs. Figure 4 shows the proposed BBs; the samples of Cardamone et al. (2009), Izotov et al. (2011), and Yang et al. (2017); and all spectroscopically classified SFGs (class_sp = 0) of the HECATE catalog. The SFRs and stellar masses plotted in the figure refer to the results of the SED fitting for the BBs of this work (Section 3.3). For the Yang et al. (2017) BBs, SFRs, and stellar masses, they are as reported in that work. For the HECATE catalog and the samples of Cardamone et al. (2009) and Izotov et al. (2011), the SFRs and stellar masses are as reported in HECATEv2.0 and the MPA-JHU catalogs, respectively.
All the aforementioned samples of dwarf galaxies show very high values of sSFR () and are placed above the main sequence compared to HECATE galaxies, which represent here the average population of SFGs in the near Universe. Moreover, BBs tend to be even more extreme compared to the Cardamone et al. (2009) GPs and Izotov et al. (2011) BDGs, with the majority being less massive and having .
@@ 3.5 Visible colors of blueberry galaxies
Figure 5 shows the rest-frame optical color-color versus diagram for this work’s BBs, the Yang et al. (2017) BBs, the Izotov et al. (2011) BDGs, and the Cardamone et al. (2009) GPs. The HECATE galaxies classified as star forming are also shown, representing the general population of SFGs in the nearby Universe. The colors plotted for both BB samples and the HECATE SFGs are based on SDSS cModelMag photometry. The colors of the BDGs and GPs are based on ModelMag photometry from SDSS DR12 (Alam et al. 2015). The photometry of all sources was -corrected using the conversions of Chilingarian et al. (2010) and Chilingarian & Zolotukhin (2012). Figure 5 shows that some BBs are redder than , which is the selection threshold adopted by this work for SDSS photometry. However, these sources were selected through Pan-STARRS photometry, which, as discussed in Section 2, can be different compared to the cModelMag plotted here (see also Table 2). Colors reported in Table 2 refer to the observed (not the -corrected) colors.
@@
The majority of this work’s BBs tend to be slightly redder in the color compared to Yang et al. (2017) BBs but extend similarly in the color. However, there are a few sources that are located at the blue-end of the color and extend to the Yang et al. (2017) BBs. Overall, this work’s BBs tend to be located in the versus diagram between the Yang et al. (2017) BBs and the Izotov et al. (2011) BDGs and Cardamone et al. (2009) GPs. This comparison also shows that the GPs and BDGs are on average intrinsically redder compared to BBs, although a small fraction of the large sample of Izotov et al. (2011) BDGs share similar blue colors with the BBs. Thus, the distinctive visible colors of BBs and GPs are not only different due to redshifting but are also intrinsically different.
3.6 Spectroscopic properties of blueberry galaxies
@@
To examine their spectroscopic properties, the BBs were cross-matched with the MPA-JHU catalog, which provides robust results based on a homogeneous spectroscopic analysis. However, this analysis could only be performed for the sources also covered by the MPA-JHU catalog, that is 21/48 galaxies. The EWs of the , H, , and H emission lines and the metallicities for the spectroscopic subsample of the proposed BBs are reported in Table 3 and are based on the MPA-JHU catalog. The distributions of the sSFRs and the EWs of the H, , and H emission lines for the GP and BB samples are presented in Figure 6. The spectroscopically classified HECATE SFGs are also shown. Emission-line fluxes and EWs for the Cardamone et al. (2009), Izotov et al. (2011), HECATE, and BBs of this work are as provided by the MPA-JHU analysis. Spectroscopic information of the Yang et al. (2017) BBs is as reported in their work, but because many of these sources were observed with poor flat-fielding of the red part of the spectrum, their were not properly estimated and are not shown in Figure 6. Table 4 presents the modes and 68% CIs of these distributions, summarizing the results shown in Figure 6.
@@
The BB galaxies of this work are in agreement and show an sSFR distribution similar to that of the BBs of Yang et al. (2017). Both samples are on the right end of the sSFR distribution, and they show the highest star-forming activity among the presented samples. While the Cardamone et al. (2009) GPs and Izotov et al. (2011) BDGs also show very high sSFRs () compared to the main population of SFGs, BBs show even higher star-formation intensity. An examination of the sample’s EW distributions revealed they had similar behavior to those of the sSFRs. This work’s BBs show an EW distribution that is similar to the Yang et al. (2017) BBs. Both samples of BBs have the highest EWs in any of the compared emission lines.
@@
The modes and CIs of the sSFRs of this work’s BBs mentioned in this table refer to the MPA-JHU estimations regarding the 21 spectroscopic sources. The modes and CIs of the sSFRs of all 48 BBs based on the SED fitting is .
@@
The distributions of log for the different samples are shown in Figure 6. The ratio of the different oxygen ions is a good tracer of the interstellar medium’s (ISM) ionization degree. The distributions of GPs and BDGs overlap to some extent with the BBs, but the modes of their distributions have lower ratios. Thus, the BBs show the most ionized ISM among the compared samples.
@@
The mass–metallicity relation of the BBs, GPs, and BDGs is presented in Figure 7. Metallicities and stellar masses of all samples, except the Yang et al. (2017) BBs, are as reported by the MPA-JHU analysis, where the metallicity estimation was based on the Charlot & Longhetti (2001) models. Stellar masses and metallicities of the Yang et al. (2017) BBs are as reported in their work. Most BBs show extremely low metallicities, with . The mode of the metallicity distribution of the Yang et al. (2017) BBs is , and for this work’s BBs, it is . For reference, we also plot in Figure 7 the mass–SFR–metallicity relation by Curti et al. (2020) provided in their Equation 2. Estimations below are an extrapolation since the Curti et al. (2020) relation was a product of an analysis with more massive galaxies compared to BBs.
@@
The highly star-forming and low-mass BBs tend to have metallicities close to what is expected from the relation accounting for their stellar mass and SFR. For instance, the mode of the SFR distribution of the plotted BBs of this work is , while the minimum and maximum SFRs are , , respectively. The upper line drawn from the Curti et al. (2020) relation refers to . Thus, the majority of the BBs are below the lines that correspond to their actual SFRs. However, the comparison shows significant scatter. Compared to BBs, GP and most BDG galaxies are more massive and show, on average, higher metallicities.
@@
Two outliers also show a higher metallicity than the rest of the BBs (BBs 2 and 33). Both of these sources show extended features rather than the typical compact morphology of the majority of the BBs, and their spectroscopic analysis refers to these regions. This suggests that these extended features are part of the same galaxy, which probably hosts older SPs and has an ISM with a higher metallicity.
@@
The comparison between the metallicity and the extinction of the aforementioned samples is shown in Figure 8. The extinction in the figure is based on the Balmer ratio and the emission-line fluxes measured by MPA-JHU for the spectroscopic subsample of the proposed BBs. The was calculated by adopting Case B recombination with (Osterbrock 1989) and the reddening law of Calzetti et al. (2000). The of sources that were found to have a Balmer ratio was set to zero. The respective extinction–metallicity relation by Kouroumpatzakis et al. (2021), which was based on a representative sample of all types of SFGs, is also drawn in the figure. Figure 8 does not show the Yang et al. (2017) sample because for some galaxies, there were problems with the red part of their spectrum. While most Cardamone et al. (2009) GPs and Izotov et al. (2011) BDGs tend to follow the relation, given the large scatter, this is not the case for the majority of BBs. Based on their metallicity, most of this work’s BBs are below the relation and show stronger extinction than expected.
@@
Table 4 also provides the modes and 68% CIs of the distributions of the quantities discussed above, separately for the morphologically compact BBs and those showing extended features. On average, compact BBs tend to form stars more actively and show slightly higher sSFRs and EWs. On the other hand, compact BBs on average seem to have higher metallicities, although their distributions overlap, and the BB with the highest metallicity (BB 2; 12 + log(O/H) = 8.6) shows extended features. Out of the 21 spectroscopically observed BBs, only six sources are reported as extended and 15 as compact. Thus, these results should be taken with caution due to the small and uneven size of these subsamples.
@@
The high ionization potential of (54.42 eV) is considered in many cases an indication of AGN activity. However, when an AGN is not responsible, such high ionization levels can only be achieved by the presence of very massive stars and recent bursts of star formation, which generate SPs with ages less than 5 Myr (e.g., Bergeron et al. 1997). The emission-line fluxes were measured by Shirazi & Brinchmann (2012) for the spectra of SDSS Data Release 7. In total, eight BBs of this work were reported by Shirazi & Brinchmann (2012) as sources with detected emission with S/N.
@@
To homogeneously estimate the fluxes for all the BBs with spectroscopic observations, we fit their SDSS spectra with the SP fitting code FADO (Gomes & Papaderos 2017). The FADO fits were based on the Bruzual & Charlot (2003) simple SP template library with metallicities and . All fits used the spectral range Å , while we adopted the recommended resolution for SDSS spectra of 2.3 Å and recessional velocity margins of 500 . The fluxes and EWs of all the emission lines estimated by FADO are provided in the online version of Table 3.
@@
A comparison between the flux estimations of our analysis and those of MPA-JHU showed excellent agreement, with , , and . The comparison of the EW estimations is , , and . Thus, FADO tends to yield slightly larger EWs on average, mainly for the Balmer lines. Out of the 21 sources with spectroscopic observations, 15 were found to have emission with , showing that the BBs are indeed dominated by extremely young SPs. In particular, these are BBs 2, 7, 9, 10, 11, 12, 14, 18, 22, 27, 30, 31, 32, 33, and 36.